ok can someone explain this to me please
if f(x)=3x+1 and g(x)= x+2
find (f-g)(x) and calculate (f-g)(5)

Answer by stanbon(75887)

if f(x)=3x+1 and g(x)= x+2
find (f-g)(x) = f(x)-g(x) = (3x+1)-(x+2) = 2x-1
--------------------------------
calculate (f-g)(5) = 2*5-1 = 9
=================================
Cheers,
Stan H.
=============